| ID | 
Date | 
Author | 
Group | 
Subject | 
Text | 
  |  
| 
  209  
 | 
25 May 2023, 01:59 | 
Jörg Hallmann | MID | Shift summary | optimization of seeding ongoing
 
 
tasks achieved during alignment:
 
 
- polarizer aligned | 
   |  | 
  208  
 | 
21 May 2023, 18:38 | 
Tokushi Sato | SPB | Shift summary | X-ray delivery
 
 
 
	9.3 keV, 800 uJ, 1-200 
pulses, stable delivery
 
 
 
Optical | 
   |  | 
  207  
 | 
19 May 2023, 18:39 | 
Tokushi Sato | SPB | Shift summary | X-ray delivery
 
 
 
	9.3 keV, 800 uJ, 1-200 
pulses, stable delivery
 
 
 
Optical | 
   |  | 
  206  
 | 
18 May 2023, 18:41 | 
Tokushi Sato | SPB | Shift summary | X-ray delivery
 
 
 
	9.3 keV, ~1 mJ, 1-30 pulses, 
stable delivery
 
 
 
Optical | 
   |  | 
  204  
 | 
14 May 2023, 05:33 | 
Tommaso Mazza | SQS | Issue | SA3 GATT has trouble in setting the pressure 
at ~0.5mbar. setting any value between 0.5 
and 0.6 will have it settle around 0.48. | 
    |  | 
  205  
 | 
13 May 2023, 22:18 | 
Maurizio Vannoni | XRO | 14.8 keV SPB 2mrad Alignment | Mirror positions and beam position on the 
imagers for 14.8 keV SPB, 2mrad
 
 
ATTENTION: the alignment was difficult | 
 6x   |  | 
  203  
 | 
12 May 2023, 23:21 | 
Romain Letrun | SPB | Shift summary | X-ray delivery
 
 
 
	9.3 keV, 1.8 mJ, 202 pulses 
most of the day.
 
	Attempted tuning | 
   |  | 
  202  
 | 
12 May 2023, 16:12 | 
Frederik Wolff-Fabris | XO | SA3 status | The SA3 EPS power limit at the Big brother 
was disabled to allow photon energy scans 
in SQS (due to Kr resonance at XGM, invalid | 
   |  | 
  201  
 | 
11 May 2023, 23:48 | 
Romain Letrun | SPB | Shift summary | X-ray delivery
 
 
 
	9.3 keV, 1.4-1.8 mJ and 
12 keV, 1.1 mJ. 202 pulses most of the day.
 
	Reduced | 
   |  | 
  200  
 | 
11 May 2023, 09:21 | 
Michaela Petrich | PRC | Attention | Dear all,
 
 
I just got the information for 
BKR that due to a failed klystron we had | 
   |  | 
  199  
 | 
11 May 2023, 05:07 | 
Konopkova Zuzana | HED | Shift summary | beam parameters: 18 keV, 1 pulse on demand, 
400 uJ
 
 
achieved: | 
   |  | 
  198  
 | 
10 May 2023, 23:23 | 
Romain Letrun | SPB | Shift summary | X-ray delivery
 
 
 
	9.3 keV, 1.6 mJ, 202 pulses 
msot of the day.
 
	Slope across | 
   |  | 
  197  
 | 
10 May 2023, 09:51 | 
karen appel | HED | Shift summary | for 9th of May - 10th of May: 
user experiment 2740 , first users with DIPOLE100-X 
long pulse laser | 
   |  | 
  196  
 | 
09 May 2023, 23:09 | 
Romain Letrun | SPB | Shift summary | X-ray delivery
 
 
 
	9.3 keV, 1.8 mJ, 202 pulses 
most fo the day. Pronounced (~30-40%) slope | 
   |  | 
  195  
 | 
08 May 2023, 09:04 | 
Jan Gruenert | PRC | Attention | Info from XRO (Maurizio Vannoni) to PRC 
concerning tunnel access 
request #152936 (Inspection and repair of | 
   |  | 
  194  
 | 
08 May 2023, 02:38 | 
Jan-Etienne Pudell | MID | Shift summary | Stable beam delivery with 23keV and 140µJ 
at 2.2, 1.1 and 0.55MHz
 
 
  | 
   |  | 
  193  
 | 
07 May 2023, 23:11 | 
Sergii, Phuong | SCS | Late shift summary 07.05.2023 | Achivments:
 
 
-Move DSSC closer to the sample.
 
 
-Measure delay scans for opposite | 
   |  | 
  192  
 | 
07 May 2023, 14:47 | 
Giuseppe Mercurio | SCS | morning shift summary | achievements
 
 
 
	exchanged He dewar and 
managed to cool down the sample again after | 
   |  | 
  191  
 | 
07 May 2023, 06:52 | 
Giuseppe Mercurio | SCS | shift summary | achievements:
 
 
 
	measured delay scans on 
16 nm samples at different magnetic fields | 
   |  | 
  190  
 | 
07 May 2023, 02:23 | 
Jan-Etienne Pudell | MID | Shift summary | We had stable x-ray delivery with 120µJ 
and up to 300 pulses.
 
 
Successful user data acquisition | 
   |   
 |